A coffee shop sold 1 cup, 7 pints, and 9 quarts of coffee. Express the number of cups of coffee sold in base two
Answer by swincher4391(1107)

1 cup = 1 cup
7 pints = 14 cups (1 pint :2 cups)
9 quarts = 36 cups (1 quart: 4 cups)
-------------------------
51 cups.
Convert 51 to base 2.
Divide 51 by 2 repeatedly and observe remainder.
51/2 = 25 r 1
25/2 = 12 r 1
12/2 = 6 r 0
6/2 = 3 r 0
3/2 = 1 r 1
1/2 = 0 r 1
Read resulting remainders from bottom to top. 110011(base 2) cups.